what are the rules to put limits on trigonometric functions?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

There are no such rules for limits.
You should know the range limit of basic trigonometric functions.
-1\leq sinx\leq 1

-1\leq cosx\leq 1
-\infty \leq tanx\leq \infty
And for the solution of trigonometric equation,the interval in the question will be provided for the solutions.